A particle starts from rest and moves around a fixed point with a constant acceleration of 2 rad.s-2.


Find its angular velocity after 2 s.


Determine the angular displacement and the number of revolutions it made after 5 s.


1
Expert's answer
2021-04-29T07:30:01-0400

Answer

Using

"\omega^2=\omega_0^2+\alpha t"

So its angular velocity after 2 s.

"\omega^2=0+2\times2\\\omega=2rad/sec."

Angular displacement

"\theta=\frac{\omega^2}{2\alpha}\\=\frac{2^2}{2\times2}=1rad"

No of revolution

=5rad.